### v1

@@ 5.2 Sensitivity Analyses
The effect of the target production rate on the comparable mass and the power consumption are summarized in Figs. 9, and 10. As can be seen, there are near-linear trends for both mass and power against the production rate. Among the analyzed range, 1 – 20 tons of oxygen per year (Fig. 9(a)), the Series Hybrid architecture is always the heaviest. While the CR architecture can be the lightest regardless of the target production for one year of operation (Fig. 9(a)), for three years of operation (Fig. 9(b)), the advantage of the WE architecture becomes clearer due to the extra methane and hydrogen required for the other architectures. When comparing the power consumption Fig. 10), the WE plant appeared to be the least power-consuming while the CR or the Series Hybrid plant requires the most power.
@@
Figure 11 shows the system mass dependence on the CR reactant, i.e., CH4 and H2, recycle rate. As can be seen in this figure, there is a linear relationship between the recycle rate and the landed mass. For one year of operation (Fig. 11(a)), the CR can be the lightest architecture. For a higher recycling rate, the difference between the WE and the Series Hybrid architecture can be almost negligible. When the operational period is three years (Fig. 11(b)), the WE can be the lightest regardless of the CR reactant recycle rate.
@@
Figure 11: System mass as a function of the CR reactant recycle rate. Target production mass is set to 10.0 t of LOX and 1.25 t of LH2.
Water ice content in the excavated regolith in a PSR also affects the comparable landed mass and the power consumption of the WE and the hybrid plants significantly (Figs. 12 and 13). Generally, as water content increases, both the landed mass and power consumption decreases. This trend affects both the lightest plant architecture and the least power-consuming plant architecture. As can be seen in Fig. 12(a), the water content has the largest impact on the mass of the WE architecture for one year operation. However, for three years’ operation, the landed mass of the Series Hybrid architecture becomes significantly lighter as the water content becomes larger. This is because as the water content increases, the Series Hybrid architecture depends more on the WE from the icy regolith rather than the CR of the dry regolith, making the system mass for dry regolith processing smaller. Since the CR architecture can be a lot heavier than the WE architecture for the three years of operation, the Series Hybrid architecture can be lighter by depending less on the CR architecture. The water content in the PSR is further considered for uncertainty analyses (Section 6.1.3).
@@ 6.1 Uncertain parameters
As Cilliers et al. [31] stated, the effect of resource uncertainty on the performance of lunar ISRU plants has not been explored enough. In Ref. [30], the impact of uncertain parameters, such as resource content and operational availability, is assessed for the hydrogen reduction process on the Moon. The results indicated the risk of oversimplification by ignoring these uncertainties.
To assess the benefits and drawbacks of each ISRU plant architecture compared in Section 5, this paper further analyzes the effects of uncertainties on the performance of the plant. Inspired by Ref. [30], Monte Carlo simulations of 5000 scenarios with different values for parameters related to the lunar environment and ISRU operations is utilized to reveal the performance of each ISRU architecture under various conditions. Table 4 summarizes the Monte Carlo parameters assumed in this study.
@@
Particle size mass fraction (wt%) $\mathcal{N}$(62.6, 8.32)
Silica content (wt%) $\mathcal{N}$(45.2, 0.932)
Water ice content (wt%) Nominal: Lognormal (ln(4), 0.32)
@@
Water yield (-) $\mathcal{N}$(0.75, 0.052)
Operational availability (days/year) $\mathcal{N}$(200, 102)
@@ 6.1.1 Particle size distribution of lunar highlands regolith
After excavation, the particle size separator removes the coarse regolith fraction (Fig. 2). In this work, to improve the conversion time of the carbothermal reduction, it is assumed that particles larger than particles larger than 150 µm are removed as mentioned in Sec. 4.3.2. A further assumption is that the lunar surface in the southern polar region is mainly covered by highlands regolith similar to Apollo 16 samples [74, 75]. Figure 15 shows the particle size distributions of 75 Apollo 16 samples reported by [76]. From these distributions, Fig. 15 is generated to depict the distribution of the mass fraction of the excavated regolith between 90 and 1000 µm. As can be observed in Fig. 15, the mass fraction is normally distributed with a mean value of 62.6 wtand a standard deviation of 8.3 wt%, i.e., $\mathcal{N}$(62.6, 8.32). Therefore, the mass fraction of excavated regolith in the southern polar region is assumed to follow this distribution.
Figure 14: Particle size distributions of 75 Apollo 16 samples obtained from [76]. The black vertical line represents 150 µm.
@@ 6.1.2 Silica content in lunar highlands regolith
As reported in [77], silica is the most dominant metal oxide in lunar highlands regolith. The Apollo 16 sample 60501 contains 45.24 wtof silica with a standard deviation of 0.93 wt%. Without further information on the silica content of the lunar southern polar regolith, we also assumed $\mathcal{N}$(45.2, 0.932) as a distribution of the silica content.
@@ 6.1.3 Water ice content in the lunar south polar region
Compared to the silica concentration, information regarding water ice content in the southern polar region of the Moon is significantly more limited. A great number of past studies have tried to estimate the icy water content in PSRs. For instance, Colaprete et al. [36] analyzed Lunar Crater Observation and Sensing Satellite’s ejecta plume and concluded that the water ice content in the impact site (Cabeus crater) was 5.6 $\pm$2.9 wt%. The authors also noted that icy water might exist heterogeneously. Hayne et al. [37] supported this heterogeneity showing some cold traps, where water ice can be thermally stable, do not have exposed water ice. This also aligns with the finding of only 3.5% of PSRs probably having surface ice by Li et al. [38]. From UV albedo values, Hayne et al. [37] concluded that 0.1–2.0 wtof water ice might be exposed in some PSRs if the ice is intimately mixed with dry regolith, whereas 1-10% of surface area would be icy if this is pure and patchy.
More recently, using the data from the Moon Mineralogy Mapper, Li et al. [38] estimated the water content in PSRs by detecting near-infrared absorption features of water ice in reflectance spectra. This study revealed that ice-bearing pixels (280 m $\times$ 280 m) may contain around 30 wtof ice intimately mixed with regolith or 20 volif pure and patchy ice exists. This is a good agreement with Zuber et al. [78], who concluded that 20% superficial thin ($\sim$1 µm) ice may exist in the Shackleton crater. As Cannon and Britt [79] argued, estimations made by Hayne et al. [37] and Li et al. [38] correspond to only surface ice (micrometers to millimeters depth), whereas the value measured from the Lunar Crater Observation and Sensing Satellite impact might be subsurface data (tenths of centimeters depth).
These past studies indicate significant heterogeneity of water ice in the southern polar region might be expected. With all the information available to this date, it is hard to conclude the water content in excavated icy-regolith. Therefore, this work tests two different scenarios: a nominal and a best-case scenario. In the nominal scenario, the large ice content reported by [38] exists only on the surface, and the actual ice content in the excavated regolith is smaller than that, which aligns with Cannon and Britt [79]. For this scenario, to avoid negative values, a log-normal distribution with a median value of 4 wtand a variance of 0.3 wt%2 is considered as the distribution of the ice content. We also test the other scenario where there indeed is more ice in the excavated regolith as reported by [38] within a reachable area from the PEL site. For this scenario, $\mathcal{N}$(30, 22) [wt%] is employed. See Fig. 16 for employed distributions as well as some reported values in past studies.
Figure 16: Water ice content.
@@ 6.1.4 Water yield
The water yield efficiency of the water extraction technology is another uncertain parameter. Kleinhenz and Paz assumed 75% efficiency [23]. Cole et al. reported 67 $\pm$ 5% of water extraction in 25 minutes of heating from highland regolith simulant [24]. From these studies, this paper assumed 75% mean efficiency with 5% of variance.
@@ 6.1.5 Operational availability
Although this work does not prescribe a detailed power system, operational uncertainty related to the power source should be considered. Assuming solar power, operational availability depends on how much sunlight the plant can receive. Mazarico et al. [39] simulated the sunlight availability of many PELs in both northern and southern polar regions (see supplementary tables 2a and 2b from Ref. [39]). According to their simulation, the best location near the ridge of the Shackleton crater can get about 202 days of continuous sunlight which aligns with the assumption made by Linne et al.[7] Therefore, a normal distribution $\mathcal{N}$(200, 82) [days per lunar year] is assumed in this paper.
@@ 6.2 Lunar ISRU performance under uncertainty
Using the varying parameters listed in Table 4, we run Monte Carlo simulations for 5000 scenarios. Figure 18 depicts the regolith excavation rate of each architecture. As can be seen in this figure, the CR and the Series Hybrid architectures require much smaller excavation rate than the WE and the Parallel Hybrid architectures. This is due to the high oxygen yield of the CR technology compared to the WE technology. The Series Hybrid architecture can potentially achieve a higher yield than the CR architecture by extracting oxygen from both regolith and water ice. However, due to the distance between a PEL and PSR, hydrogen generated in the PEL is not recycled to Reaction 3 in this architecture. This limitation leads to the yield of this Series Hybrid architecture being lower than the CR architecture. The uncertainty in the water ice content makes the required regolith excavation rates for the WE and the Parallel Hybrid architectures unclear. For instance, for the WE architecture, 82 kg/h of regolith excavation is required to produce 10 t of liquid oxygen and 1.25 t of liquid hydrogen with a 50% confidence. However, to achieve a 95% confidence, the required excavation rate becomes 130 kg/h. According to [62], RASSOR is designed to excavate at least 2.7 t of regolith per day (112.5 kg/h). Therefore, depending on the condition, the WE architecture has a risk of not meeting the excavation mass target unless a redundant excavator is considered.
@@
Figure 18 shows the histograms of power consumption of each architecture with varying parameters. As can be seen, the average required power of the hybrid architectures and the CR architecture does not show significant differences. Although the average power consumption of the WE architecture is lower than the other architectures, due to the larger uncertainty in the water ice content, the histogram of the WE architecture shows the widest variety in its power consumption. The non-normality observed in this figure is due to the non-continuity of some design parameters. For instance, the number of molten zones for the CR process and the number of cryocoolers to liquefy oxygen and hydrogen must be integers, making several peaks in the histograms.
Figure 17: Excavation rate to meet target production, 10 t of LOX and 1.25 t of LH2.
Figure 18: Power consumption to produce 10 t of LOX and 1.25 t of LH2.
@@
Figures 20 and 20 are the results of the Monte Carlo simulations with the best-case water content (i.e., $\mathcal{N}$(30, 22)). With this abundance of water ice in the lunar southern pole, the the Series Hybrid architecture outperforms the other architectures in terms of the regolith excavation rate (Fig. 20). For power consumption (Fig. 20), the WE plant is expected to require the least. The similar non-normality seen in Fig. 18 is observable in Fig. 20, too.
@@
Figure 19: Excavation rate to meet target production, 10 t of LOX and 1.25 t of LH2. The best case for water ice content distribution: $\mathcal{N}$(30, 22).
@@
Figure 20: Power consumption to produce 10 t of LOX and 1.25 t of LH2. The best case for water ice content distribution: $\mathcal{N}$(30, 22).
@@
Finally, the required comparable mass including the additional hydrogen and methane for one year’s operation are summarized in Fig. 21. For both water ice content distributions, the CR plant mass variance is significantly smaller than the others. In terms of the expected system mass, the lightest architecture is the CR while the heaviest is the Series Hybrid architecture. For the best-case water ice content (Fig. 21(b)), the heavier ends of the required landed mass distributions of the WE and the Series Hybrid architectures show similar values. This indicates the required landed mass of these architectures in the worst case could be almost the same.
(a) Water ice content: Log-normal (ln(4), 0.32).
(b) Water ice content: $\mathcal{N}$(30, 22).
@@ 7.1 Summary of key findings
For power consumption, our study shows that the WE architecture can be often the least power-consuming design (Figs. 8, 10, and 13). It should be noted that the CR technology in this paper requires direct thermal power instead of electrical power. This can be particularly challenging for the Series Hybrid architecture since the CR process of this hybrid architecture is conducted in a PSR. This requires a mirror system suggested in [15], indicating more landed mass is required compared to the other architectures.
From the Monte Carlo studies (Sec. 6), the WE architecture turns out to require the largest excavation rate range, highlighting the challenge of engineers to optimize its design. This can also be seen in Fig. 21 as the widest range of the required system mass for the WE plant. Table 5 summarizes the required mass with a 50% confidence to meet the target production rate and those with a 95% confidence. To improve the robustness of each plant design, 95% confidence is more desirable. The WE architecture requires 750 kg of additional mass to achieve this confidence while the CR architecture only requires about 75 kg.
@@
(a) Water ice content: Log-normal (ln(4), 0.32).
@@
the target production CR WE Parallel Series
50% 4,406 6,064 5,319 6,762
95% 4,483 6,828 5,707 7,368
(b) Water ice content: $\mathcal{N}$(30, 22).
@@
the target production CR WE Parallel Series
50% 4,406 5,754 5,072 6,019
95% 4,483 6,493 5,389 7,673

### v2

@@ 5.2 Sensitivity Analyses
As mentioned, the CR process requires additional H2 and CH4; therefore, the landed mass of any architecture with the CR process changes depending on the operational period. Figure 10 shows the change in the landed mass of each architecture in response to varying operational periods. Each architecture is designed to produce 10.0 t of LOX and 1.25 t of LH2 annually. As can be seen in this figure, the order of architectures except for the SH is reversed between one year and two years’ operation regarding their landed mass. For six years or longer operation, the CR architecture requires the most landed mass.
Figure 10: System mass as a function of operational period. The target production mass and the CR reactant recycle rate are set to 10.0 t of LOX and 1.25 t of LH2, and 91%, respectively
The effect of the target production rate on the comparable mass and the power consumption are summarized in Figs. 11, and 12. As can be seen, there are near-linear trends for both mass and power against the production rate. Among the analyzed range, 1 – 20 t of oxygen per year (Fig. 11(a)), the SH architecture is always the heaviest. While the CR architecture can be the lightest regardless of the target production for one year of operation (Fig. 11(a)), for three years of operation (Fig. 11(b)), the advantage of the WE architecture becomes clearer due to the extra methane and hydrogen required for the other architectures. When comparing the power consumption (Fig. 12), the CR plant appeared to be the least power-consuming while the WE plant requires the most power for the majority of the studied range. Compared to the landed mass, the variance in the power consumption among all architectures remains relatively small.
(a) One-year operation.
(b) Three-years operation.
@@
Figure 13 shows the system mass dependence on the CR reactant, i.e., CH4 and H2, recycle rate. As can be seen in this figure, there is a linear relationship between the recycle rate and the landed mass. For one year of operation (Fig. 13(a)), the CR can be the lightest architecture. For a higher recycling rate, the difference between the WE and the SH architecture can be almost negligible. When the operational period is three years (Fig. 13(b)), the WE can be the lightest regardless of the CR reactant recycle rate.
@@
Figure 13: System mass as a function of the CR reactant recycle rate. The target production mass is set to 10.0 t of LOX and 1.25 t of LH2.
Water ice content in the excavated regolith in a PSR also affects the comparable landed mass and the power consumption of the WE and the hybrid plants significantly (Figs. 14 and 15). Generally, as water content increases, both the landed mass and power consumption decreases. This trend affects the heaviest plant architecture, as well as the most and the least power-consuming plant architectures. As can be seen in Fig. 14(a), the water content has the largest impact on the mass of the WE architecture for one year operation. However, for three years’ operation, the landed mass of the SH architecture becomes significantly lighter as the water content becomes larger. This is because as the water content increases, the SH architecture depends more on the WE from the icy regolith rather than the CR of the dry regolith, making the system mass for dry regolith processing smaller. Since the CR architecture can be a lot heavier than the WE architecture for the three years of operation, the SH architecture can be lighter by depending less on the CR architecture.
Regarding power requirements, the WE architecture is affected the most by the water content due to its sole reliance on the WE process. As a result, the WE architecture consumes the most power when the water content is smaller than or equal to 4 wtwhile it becomes the least power-consuming architecture when the water content is 6 wtor higher. In contrast, the SH architecture requires relatively low power across the analyzed range. This is because the SH architecture significantly depends on the CR process more when the water content is low, and gradually shifts its reliance to the WE process as water content increases. The water content in the PSR is further considered for uncertainty analyses (Section 6.1.3).
@@ 6.1 Uncertain parameters
As indicated by the study by Cilliers et al. [30], there exists the risk of underperformance when various uncertainties, such as in resource content, are not considered. To assess the benefits and drawbacks of each ISRU plant architecture compared in Section 5, this paper further analyzes the effects of uncertainties on the performance of the plant. Inspired by the aforementioned study by Cilliers et al., Monte Carlo simulations of 5000 scenarios with different values for parameters related to the lunar environment and ISRU operations is utilized to reveal the performance of each ISRU architecture under various conditions. Table 5 summarizes the Monte Carlo parameters assumed in this study.
@@
Particle size mass fraction [wt%] $\mathcal{N}$(62.6, 8.32)
Silica content [wt%] $\mathcal{N}$(45.2, 0.932)
Water ice content [wt%] Nominal: Lognormal (ln(4), 0.32)
@@
Water yield [-] $\mathcal{N}$(0.75, 0.052)
Operational availability [days/year] $\mathcal{N}$(200, 102)
@@ 6.1.1 Particle size distribution of lunar highlands regolith
After excavation, the particle size separator removes the coarse regolith fraction (Fig. 3). In this work, to improve the conversion time of the CR, it is assumed that particles larger than particles larger than 150 µm are removed as mentioned in Sec. 4.3.2. A further assumption is that the lunar surface in the southern polar region is mainly covered by highlands regolith similar to Apollo 16 samples [87, 88]. Figure 16 shows the particle size distributions of 75 Apollo 16 samples reported by Graf [89]. From these distributions, Fig. 17 is generated to depict the distribution of the mass fraction of the excavated regolith between 90 and 1000 µm. As can be observed in Fig. 17, the mass fraction is normally distributed with a mean value of 62.6 wtand a standard deviation of 8.3 wt%, i.e., $\mathcal{N}$(62.6, 8.32). Therefore, the mass fraction of excavated regolith in the southern polar region is assumed to follow this distribution.
Figure 16: Particle size distributions of 75 Apollo 16 samples obtained from Graf [89]. The black vertical line represents 150 µm.
@@ 6.1.2 Silica content in lunar highlands regolith
As reported by Ridley et al. [90], silica is the most dominant metal oxide in lunar highlands regolith. The Apollo 16 sample 60501 contains 45.24 wtof silica with a standard deviation of 0.93 wt%. Without further information on the silica content of the lunar southern polar regolith, we also assumed $\mathcal{N}$(45.2, 0.932) as a distribution of the silica content.
@@ 6.1.3 Water ice content in the lunar south polar region
Compared to the silica concentration, information regarding water ice content in the southern polar region of the Moon is significantly more limited. A great number of past studies have tried to estimate the icy water content in PSRs. For instance, Colaprete et al. [36] analyzed Lunar Crater Observation and Sensing Satellite’s ejecta plume and concluded that the water ice content in the impact site (Cabeus crater) was 5.6 $\pm$2.9 wt%. The authors also noted that icy water might exist heterogeneously. Hayne et al. [37] supported this heterogeneity showing some cold traps, where water ice can be thermally stable, do not have exposed water ice. This also aligns with the finding of only 3.5% of PSRs probably having surface ice by Li et al. [38]. From UV albedo values, Hayne et al. concluded that 0.1–2.0 wtof water ice might be exposed in some PSRs if the ice is intimately mixed with dry regolith, whereas 1-10% of surface area would be icy if this is pure and patchy.
More recently, using the data from the Moon Mineralogy Mapper, Li et al. [38] estimated the water content in PSRs by detecting near-infrared absorption features of water ice in reflectance spectra. This study revealed that ice-bearing pixels (280 m $\times$ 280 m) may contain around 30 wtof ice intimately mixed with regolith or 20 volif pure and patchy ice exists. This is a good agreement with Zuber et al. [91], who concluded that 20% superficial thin ($\sim$1 µm) ice may exist in the Shackleton crater. As Cannon and Britt [92] argued, estimations made by Hayne et al. [37] and Li et al. correspond to only surface ice (micrometers to millimeters depth), whereas the value measured from the Lunar Crater Observation and Sensing Satellite impact might be subsurface data (tenths of centimeters depth).
These past studies indicate significant heterogeneity of water ice in the southern polar region might be expected. With all the information available to this date, it is hard to conclude the water content in excavated icy-regolith. Therefore, this work tests two different scenarios: a nominal and a best-case scenario. In the nominal scenario, the large ice content reported by Li et al. [38] exists only on the surface, and the actual ice content in the excavated regolith is smaller than that, which aligns with Cannon and Britt [92]. For this scenario, to avoid negative values, a log-normal distribution with a median value of 4 wtand a variance of 0.3 wt%2 is considered as the distribution of the ice content. We also test the other scenario where there indeed is more ice in the excavated regolith as reported by Li et al. within a reachable area from the PEL site. For this scenario, $\mathcal{N}$(30, 22) [wt%] is employed. See Fig. 18 for employed distributions as well as some reported values in past studies.
Figure 18: Water ice content. Values from past studies [36, 37, 38] and the distributions considered in this paper.
@@ 6.1.4 Water yield
The water yield efficiency of the WE technology is another uncertain parameter. Kleinhenz and Paz assumed 75% efficiency [23]. Cole et al. reported 67 $\pm$ 5% of WE in 25 minutes of heating from highland regolith simulant [24]. From these studies, this paper assumed 75% mean efficiency with 5% of variance.
@@ 6.1.5 Operational availability
Although this work does not prescribe a detailed power system, operational uncertainty related to the power source should be considered. Assuming solar power, operational availability depends on how much sunlight the plant can receive. Mazarico et al. [39] simulated the sunlight availability of many PELs in both northern and southern polar regions (see supplementary tables 2a and 2b). According to their simulation, the best location near the ridge of the Shackleton crater can get about 202 days of continuous sunlight which aligns with the assumption made by Linne et al. [7] Therefore, a normal distribution $\mathcal{N}$(200, 82) [days per year] is assumed in this paper.
@@ 6.2 Lunar ISRU performance under uncertainty
Using the varying parameters listed in Table 5, we run Monte Carlo simulations for 5000 scenarios. Each scenario selects a different parameter set following distributions listed in Table 5. Figure 20 depicts the regolith excavation rate of each architecture. As can be seen in this figure, the CR and the SH architectures require much smaller excavation rate than the WE and the PH architectures. This is due to the high oxygen yield of the CR technology compared to the WE technology. The SH architecture can potentially achieve a even higher yield than the CR architecture by extracting oxygen from both regolith and water ice. However, due to the distance between a PEL and PSR, hydrogen generated in the PEL is not recycled to the methanation process (Equation 3) in this architecture. This limitation leads to the yield of the SH and CR architecture almost the same. The uncertainty in the water ice content makes the required regolith excavation rates for the WE and the PH architectures unclear. For instance, for the WE architecture, 79 kg/h of regolith excavation is required to produce 10 t of liquid oxygen and 1.25 t of liquid hydrogen with a 50% confidence. However, to achieve a 95% confidence, the required excavation rate becomes 132 kg/h. According to Mueller et al. [52], RASSOR is designed to excavate at least 2.7 t of regolith per day (112.5 kg/h). Therefore, depending on the condition, the WE architecture has a risk of not meeting the excavation mass target unless a redundant excavator is considered.
@@
Figure 20 shows the histograms of the power consumption of each architecture with varying parameters. As can be seen, the average required power of the WE architecture is the highest and that of the CR is the lowest. The average power values of the SH and the PH are between those of the CR and WE. Although the average power consumption of the WE architecture is higher than the other architectures, due to the larger uncertainty in the water ice content, the histogram of the WE architecture shows the widest variety in its power consumption. The lower bound of power consumption for both the CR and WE architectures is similar. Specifically, at the 5% confidence level, the required power for the WE architecture is approximately 52 kW—only about 1 kW higher than that of the CR architecture. The non-normality observed in this figure is due to the non-continuity of some design parameters. For instance, the number of molten zones for the CR process and the number of cryocoolers to liquefy oxygen and hydrogen must be integers, making several peaks in the histograms.
[…]
@@
Figures 22 and 22 are the results of the Monte Carlo simulations with the best-case water content (i.e., $\mathcal{N}$(30, 22) [wt%]). With this abundance of water ice in the lunar southern pole, the the SH architecture significantly outperforms the other architectures in terms of the regolith excavation rate (Fig. 22). For power consumption (Fig. 22), the WE plant is expected to require the least followed by the SH. The similar non-normality seen in Fig. 20 is observable in Fig. 22, too.
@@
Figure 21: Excavation rate to meet target production, 10 t of LOX and 1.25 t of LH2. The best case for water ice content distribution: $\mathcal{N}$(30, 22) [wt%].
@@
Figure 22: Power consumption to produce 10 t of LOX and 1.25 t of LH2. The best case for water ice content distribution: $\mathcal{N}$(30, 22) [wt%].
@@
Finally, the required comparable mass including the additional hydrogen and methane for one year’s operation are summarized in Fig. 23. For both water ice content distributions, the CR plant mass variance is significantly smaller than the others. In terms of the expected system mass, the lightest architecture is the CR while the heaviest is the SH architecture.
(a) Water ice content: Log-normal (ln(4), 0.32) [wt%].
(b) Water ice content: $\mathcal{N}$(30, 22) [wt%].
@@ 7.1 Summary of key findings
For power consumption, while the CR architecture requires the least power in the baseline case (Fig. 9), it is revealed that the promising architectures change drastically depending on the water content (Figs. 15, 20, and 22). It should be noted that the CR technology in this paper requires direct thermal power instead of electrical power. This can be particularly challenging for the SH architecture since the CR process of this hybrid architecture is conducted in a PSR. This requires a mirror system suggested by Sowers and Dreyer [15], indicating more landed mass is required compared to the other architectures.
From the Monte Carlo studies (Sec. 6), the WE architecture turns out to require the largest excavation rate range, highlighting the challenge of engineers to optimize its design. This can also be seen in Fig. 23 as the widest range of the required system mass for the WE plant. Table 6 summarizes the required mass with a 50% confidence to meet the target production rate and those with a 95% confidence. To improve the robustness of each plant design, 95% confidence is more desirable. With the nominal assumption of water content, the WE architecture requires 1,017 kg of additional mass to achieve this confidence while the CR architecture only requires about 185 kg.
@@
(a) Water ice content: Log-normal (ln(4), 0.32) [wt%].
@@
the target production CR WE PH SH
50% 5,414 7,428 6,429 8,202
95% 5,600 8,445 6,897 8,956
(b) Water ice content: $\mathcal{N}$(30, 22) [wt%].
@@
the target production CR WE PH SH
50% 5,415 6,932 6,159 7,506
95% 5,596 7,785 6,574 8,347